A boy shoves his stuffed toy zebra down a frictionless chute. It starts at a height of 1.89 m above the bottom of the chute with an initial speed of 1.37 m/s . The toy animal emerges horizontally from the bottom of the chute and continues sliding along a horizontal surface with a coefficient of kinetic friction of 0.293 . How far from the bottom of the chute does the toy zebra come to rest? Assume ?=9.81 m/s2 .
Answer #1

  initial energy E = PE + KE

= M * 9.81m/s² * 1.89m + ½ * M * (1.37m/s)² = M * 19.479m²/s²


friction work done = M * 19.479m²/s²

= µmgd = 0.293 * M * 9.81m/s² * d


d = 6.776 m